OpenStudy (anonymous):

give exact and approximate solutions to the tird decimal places x^2+16x+64=9 x=?

OpenStudy (anonymous):

x=-11,-5?

OpenStudy (anonymous):

x^2 + 16x + 55 = 0 you can factor this (x+ 11)(x+5) = 0 x = -11, -5 yup, you're right :)

OpenStudy (anonymous):

ty :)